OpenStudy (anonymous):

The width of a rectangle is 3 feet, and the diagonal is 5 feet. What is the area of the rectangle?

OpenStudy (anonymous):

789349 FEET ITS MATH

OpenStudy (anonymous):

find the lenght first l^2=5^2 -3^2=16 or l=4 arear =4*3 =12

OpenStudy (noelgreco):

Sounds like the simplest Pythagorean triple.